A satellite revolves in a circular orbit at a height of 200km from the surface of Earth. If the period of revolution of satellite is 90 minutes, G=6.66*10^-11 and mean radius is 6*10^6 m , calculate the average density of Earth

Grade:11

1 Answers

Khimraj
6 years ago
Time period of revolution of satellite T = 2π(r3/G*mE)1/2 
So T2 = 4π2(r3/G*mE)
(G*mE)/r3 = 4π2/T2
So (4π/3)G*ρE =  4π2/T2
Then ρ= 3π/(T2*G) = 3π/[(90*60)2 * 6.66*10-11]
SO ρ= 4853 Kg/m3
Hope it clears. If still you have doubt you can ask me.
If you like answer then please approve it.